# The Failed Migration of Academic Twitter

### v2

@@ 4.2.2. Internal Networks aggregated by Server
Figure 3. internal field network (size: number of scholars in each field; edge weights: frequency of connections; edge color: same as the source).
4.2.3. Case study: Mastodon Community of Information Security
We also examined a discipline-specific server, unique to decentralized social platforms, and its user retention rate using a case study of the Information Security discipline. We chose this discipline based on our field-level engagement analysis shown in Figure A1 in Appendix, in the Appendix, which indicates that scholars in physical sciences and mathematics were among the most active on Mastodon. To further understand the impact of server instance choice on community formation, we visualized their internal network, highlighting the top five servers where these scholars are based. Figure A2 shows the internal follower-followee network for physical science and mathematics field, highlighting the top five servers from a total of 33 servers that these scholars use. Interestingly, scholars from Information Security formed a strongly connected community on a custom discipline specific server, unlike other disciplines that used general servers such as mastodon.social or fediscience.org.
@@ 4.2.3. Case study: Mastodon Community of Information Security
To explore how a discipline specific server could influence user retention and engagement, we isolated the internal network of accounts within Information Security discipline. Figure 5 shows the internal network of the discipline, where most scholars congregated on a single server: infosec.exchange (highlighted in red). We compare the retention roles of scholars from the discipline with those from a discipline where users were concentrated in a general-purpose server and other sub-fields of similar size. We see a significantly high percentage of persistent users for information security scholars (56.8%) in comparison to other comparable disciplines. This suggests that establishing field- or discipline-specific communities or servers could foster a stronger sense of belonging, which may help retain scholars who have migrated, reducing the likelihood of their return to Twitter or seeking alternative platforms.
Figure 4. internal server network (removed isolated servers, red are servers with more than 100 scholars).
Figure 5. internal follower-followee network for Information Security community (removed isolates).
Category Info. Sec. Education Communication & Media Studies
Count Percent Count Percent Count Percent
One-Time 23 8.7 52 21.3 41 15.5
Short-Term 31 11.7 64 26.2 53 20.0
Long-Term 60 22.7 59 24.2 81 30.6
Persistent 150 56.8 69 28.3 90 34.0
Total 264 100.0 244 100.0 265 100.0
@@
Table 3. User retention breakdown for Information Security scholars compared to the sub-fields that are similar in size.
4.3. Twitter-Mastodon Cross-platform Analysis
@@
Since our attrition analysis showed that only a few users remained active on Mastodon, we conducted a cross-platform analysis to analyze their account state on Twitter. Approximately 50% of the scholars voluntarily provided their Twitter usernames, which were accurate and traceable to their Twitter profile pages. Using the Official Twitter Basic API, we collected the profile information of these accounts. Despite some account deletions, we successfully retrieved 3,131 Twitter accounts that could be matched to Mastodon accounts. Next, we retrieved the most recent posting times of these Twitter accounts to assess their activity status.
@@
Locked: Scholar’s account is locked down, meaning their posts are private and cannot be accessed.
Inactive: Scholar’s latest tweet was posted before December 31st, 2022, and their posts were made before this date, indicating no recent activity since the start of the year.
Semi-active: Scholar’s latest tweet was posted between December 31st, 2022 and November 1, 2023. This indicates some activity during the year but no recent posts after November, 2023, the last timestamp of our Mastodon tracking.
Active: Scholar’s latest tweet was posted after November 1st, 2023, indicating continuous activity and engagement on the platform.
Using the above categorization, we calculated the count and connection statistics for each category, and all statistics are presented in Table 4 (left). We observe that 79.72% of scholars are still active on Twitter, indicating no signs of migration or intention to fully migrate. These active Twitter users have accumulated a much higher number of followers, which make it more difficult to give up the platform. Some active users still make their Mastodon information accessible to the community. For instance, one scholar noted:
”Hi ya’ll for those moving over to #Mastodon you can currently find me ¡Scholar’s Mastodon Account¿.”
@@
”I’m leaving Twitter but I’ve been advised not to delete my account in order to avoid possible misuse - so I’ll leave it here as a placeholder. See you at Mastodon!”
@@
Figure 6. Connection metrics mapping between mastodon and Twitter.
Count Following Count Follower Count One-time Short-term Long-term Persistent
Active 2496 Mean:1989.07, std:3087.68 Mean:13208.76 , std:108495.54 374 578 684 860
Semi-active 220 Mean:1137.56 , std:1263.87 Mean:4467.28, std:19152.90 39 42 43 96
Inactive 91 Mean:814.56 , std:1062.39 Mean:2401.46 , std:5203.50 12 19 21 39
Locked 324 Mean:1859.72 , std:2300.91 Mean:6347.10 , std:31583.98 51 51 82 140
Table 4. Engagement statistics of 3,131 scholars and mapping of Twitter (Active, Inactive, Protected) and Mastodon (One Time, Short Term, Long Term, Persistent) activity levels.
@@
We cross-mapped Twitter activity levels with Mastodon role categories, as shown in Table 4 (right). Our analysis reveals that 49.31% of the sampled scholars are relatively active on both Mastodon (long-term or persistent users) and Twitter (active users). Among those who are inactive on Twitter or have locked their accounts, 67.95% are long-term or persistent users of Mastodon. However, this group of dedicated Mastodon users represents only 9.02% of the total scholars included in the cross-platform analysis. This suggests that while there is a core group of loyal users who have fully migrated to Mastodon, they make up a relatively small portion of the overall academic community.
@@
We also plotted the pairwise follower and following counts for the Mastodon and Twitter accounts of the same 3,131 scholars, as shown in Figure 6. The difference in connection scale is clear: while scholars who migrated to Mastodon made efforts to establish a new community, it is nowhere near the scale of connections they built over the years on Twitter. Additionally, we plotted the Twitter account creation date against the total number of tweets, displayed in Figure A3. The color of the points represents the total number of likes received by each account. A great proportion of scholars created their Twitter accounts between 2008 and 2012, indicating that the deep, long-standing networks, high levels of engagement, and rich posting histories developed on Twitter are not easily replicable or discarded. This entrenched presence highlights the difficulty of transitioning to a new platform like Mastodon without sacrificing years of accumulated influence and community ties.
4.3.1. Alternatives to Mastodon
@@
In addition to Mastodon, other centralized alternatives to Twitter have emerged in recent years, and many academics explored these venues. Unlike Mastodon, these platforms provided familiar interfaces and ease of transitioning. Those present on these alternative platforms often added their account information in their Twitter account descriptions to make it easier for others to discover and connect with them on these new platforms. To understand the extent to which academics were exploring the alternatives,we analyzed their usernames and user descriptions for mentions of these platforms. Specifically, we focused on mentions of Bluesky [^5] https://bsky.app/ and Threads [^6] https://www.threads.net/ using keyword matching as well as user name pattern matching. For Mastodon and Bluesky, we also included scholars who use decentralized platform-style usernames as their Twitter usernames (e.g., ¡username¿@¡servername¿).
@@
Figure 7. Mentions of Platforms in Username and Profile Description.
4.4. Analysis of Tweets on Mastodon Migration
To further investigate the conversations surrounding Mastodon among academics, we leverage data collected from Twitter reflecting conversations about academics on Mastodon (see Section 3). We plot the daily frequency of these tweets and engagement counts over time aggregated by week in Figure 8. Discussions about Mastodon in the academic context experienced a noticeable spike during the first two months of the movement. After the migration period, although conversations were less frequent, they received more attention (likes), indicating that the audience and connections of academics concerned about Mastodon have largely remained active on Twitter.
Conversations reflect difficulties and concerns during the transition. For example, Mastodon struggled to accommodate the abrupt influx of users. On the peak day of discussion, November 18, 2022:
”I have had a Mastodon presence for several months but haven’t used it. I still have a vain hope Twitter will sort itself out, but if not, I am ¡scholar’s Mastodon account¿.”
”As you know, I’m over on Mastodon too. Personally I don’t think that I will leave Twitter until and unless Elon finally and truly kills it, but I know that we will connect over on mastodon if he does.”
@@ 4.4. Analysis of Tweets on Mastodon Migration
Figure 8. Dynamic frequency of Twitter conversations related to academics migrating to Mastodon and corresponding engagement metrics (connection and like counts), aggregated by week.
@@
”Vibe-wise, mastodon has *a ton* of different servers already, so there’s more diversity. but the *big* servers that most ppl are on are… kinda academic? kinda techy? very… not fun? it feels like linkedin or facebook groups to a lot of people.”
”My humanities folks have largely migrated to mastodon, but my science and public health feeds are still here on Twitter. More’s the pity.”
Despite these challenges, ongoing efforts to promote Mastodon indicate a persistent belief in its potential. For instance, in June 2023, one scholar posted,
”Reasons to join us in the fediverse. 500 character limit. free edit button. Lots more #openaccess resources for #teachers, researchers, science communicators, lifelong learners. #STEM #education #openscience #K12 #HigherEd”
Amongst the discussions, we again observe concurrent mentions of other platforms such as Bluesky and Threads. These have diluted the migration as users explore various platforms for different purposes. For instance,
”I’m curious re categories of people I follow and where they go. IT people seem to have mostly gone to mastodon; librarians ditto; academic leftists stayed here longest, now going to bluesky (prefer commercial sites?); archaeologists have vanished for me completely (threads?).”
”Hey if anyone’s still here on science Twitter and wondering where a lot of people have gone, they’re trying out bluesky and liking it. This time it’s real (not like mastodon). #academic”
From a qualitative perspective, these conversations reflect both optimism in Mastodon’s potential and the practical concerns academics face during the migration.
5. Discussion and Conclusion
@@
Results from longitudinal analysis of Mastodon user data showed that the initial surge in sign-ups did not translate into sustained long-term user engagement. Most recent user activity on Twitter reveals that many academics in our dataset returned to using Twitter actively. Further analysis of engagement metrics reveals that the level of established history, as well as the strong communities established on Twitter, with some over a decade, proved too significant to overcome. Without a complete migration of user history, and established community, professionals will naturally gravitate back to familiar platforms where their communities are already established through years.
Several factors likely contributed to the failure of Mastodon to retain users. First, its decentralized structure comes with a steep learning curve compared to Twitter. Furthermore, competition from more user-friendly platforms like Bluesky and Threads potentially drew users away from Mastodon. Additionally, those who continued to post consistently on Mastodon, did not receive same level of engagement as on Twitter. While some users expressed doubts, certain groups of academics stayed committed to Mastodon, making this migration less of a straightforward failure and more of a complex transition.
One significant limitation of this study is the inconsistency in data collection timing across different platforms. This discrepancy could potentially introduce biases in the analysis, as activity levels of scholars might vary depending on when the data was captured. Additionally, incorrect or incomplete social media handles provided by the scholars led to gaps in data that we could not manually resolve due to the scale of the dataset. Moreover, the changes to platform access (such as Twitter API rate limitations) restricted our ability to gather sufficient data throughout the study. However, we believe our findings from the remaining data is substantial and provides valuable insights for the cross platform analyses.

### v3

@@ Introduction
Figure 2: Internal field network (size: number of scholars in each field; edge weights: frequency of connections; edge color: same as the source; self-loops are omitted for better visualization.)
Figure 3: (Top) Server network constructed from internal follower–followee relationships. Isolated servers are removed for clarity. Node size is scaled by the number of scholars within each server. Servers with more than 100 scholars are highlighted in red. (Bottom) Internal comment interaction network among servers with more than 100 scholars. Node size is scaled by the total number of interactions for each server, and edge width is scaled by the number of interactions between servers. Internal boost, favorite, and mention networks are reported in Figure A2 in the Appendix.
Internal networks aggregated by server
Prior studies have shown that the Mastodon’s server instance based structure makes it easier for users to discover and connect with others within the same instance than outside instance. This leads to stronger within instance ties compared to weaker links across instances (Zignani et al. 2019). Therefore, we analyzed the follower-followee network aggregated by Mastodon instances and the corresponding interaction networks among top servers, shown in Figure 3.
The academic users were spread across a total of 342 servers, with 9 server instances being isolates without any connecting edges to other instances, which were removed for clarity. Table A3 in the Appendix shows network centrality metrics for each server and the distribution of the users across different servers. It also includes the official description of the server instances, indicating whether they are customized to a specific group of scholars. During the migration, different disciplines used different strategies: some used general-purpose servers like mastodon.social which host a wide variety of content beyond just academic content, while others took a coordinated approach on setting up their field specific server instances such as infosec.exchange and journa.host. Figure 3 (Top) shows the follower-followee connections across different servers showing certain general purpose instances such as mastodon.social are popular among the tracked academic users (containing 21.3% of the total) than other general purpose servers like mas.to. We observe a central hub pattern in which the majority of follower–followee connections involve the mastodon.social server and other servers, forming a radiating structure. To examine interactions beyond follower relationships, we further plot interaction networks among top servers with more than 100 scholars (highlighted in red in Figure 3 (Top)). Figure 3 (Bottom) shows the internal comment interaction network (boost, favorite, and mention networks are presented in Figure A2 in the Appendix). Node size is scaled by the total number of interactions for each server, and edge width is scaled by the number of interactions between server pairs. We find that even when servers are not directly connected through follower–followee relationships, they are connected through interactions such as comments, mentions, and boosts. This motivates the inclusion of both within-server and inter-server interaction in subsequent modeling, as these interaction-based connections capture federated engagement that is not visible in the follower network alone. In particular, interaction networks reveal cross-server communication pathways that operate independently of stable follower ties.
@@
Comparison Across Communities. To explore how a discipline specific server could influence user retention and engagement, we isolated the internal network of accounts within Information Security. Figure A3 shows the internal network of the discipline, where most scholars congregated on a single server: infosec.exchange (highlighted in red). We compare the retention roles of scholars from the discipline with those from a discipline where users were concentrated in a general-purpose server and other sub-fields of similar size (Education and Communication & Media Studies), as shown in Figure 4. We see a significantly higher percentage of persistent users amongst information security scholars (40.9%) in comparison to other comparable disciplines. This suggests that establishing field- or discipline-specific communities or servers could foster a stronger sense of belonging, which may help retain scholars who have migrated, reducing the likelihood of their return to Twitter or seeking alternative platforms.
@@
Figure 4: User retention breakdown for Information Security scholars compared to the sub-fields that are similar in size (Info. Sec.: n=264; Comm. & Media Studies: n=265; Education: n=244.)
Cross-Platform Patterns and Usage Shifts
@@
Since our attrition analysis categorized 1845 users as active on Mastodon, we conducted a cross-platform analysis to analyze their account state on Twitter. Approximately 50% of the scholars voluntarily provided their Twitter usernames, which were accurate and traceable to their Twitter profile pages. Using the Official Twitter Basic API, we collected the profile information of these accounts. Despite some account deletions, we successfully retrieved 3,131 Twitter accounts that could be matched to Mastodon accounts. Next, we retrieved the most recent posting times of these Twitter accounts to assess their activity status.
Count Following Count Follower Count
Active 2496 $\mu$:1989.07, $\sigma$:3087.68 $\mu$:13208.76 , $\sigma$:108495.54
Semi-active 220 $\mu$:1137.56 , $\sigma$:1263.87 $\mu$:4467.28, $\sigma$:19152.90
Inactive 91 $\mu$:814.56 , $\sigma$:1062.39 $\mu$:2401.46 , $\sigma$:5203.50
Locked 324 $\mu$:1859.72 , $\sigma$:2300.91 $\mu$:6347.10 , $\sigma$:31583.98
Table 2: Engagement statistics of 3,131 scholars (number of users per category with mean and standard deviation of following and follower counts).
@@
Locked. Scholar’s account is locked down, meaning their posts are private and cannot be accessed.
Inactive. Scholar’s latest tweet was posted before December 31st, 2022, and their posts were made before this date, indicating no recent activity since the start of the year.
Semi-active. Scholar’s latest tweet was posted between December 31st, 2022 and November 1, 2023. This indicates some activity during the year but no recent posts after November, 2023, the end of our Mastodon tracking.
Active. Scholar’s latest tweet was posted after November 1st, 2023, indicating continuous activity and engagement on the platform.
Using the above categorization, we calculated the count and connection statistics for each category, and all statistics are presented in Table 2. We observe that 79.72% of scholars are still active on Twitter, indicating no signs of migration or intention to fully migrate. These active Twitter users have accumulated approximately 10 times more followers, which make it more difficult to give up the platform. Some active users still make their Mastodon information accessible to the community. For instance, one scholar noted:
“Hi ya’ll for those moving over to #Mastodon you can currently find me <<Scholar’s Mastodon Account>>.”
@@
“I’m leaving Twitter but I’ve been advised not to delete my account in order to avoid possible misuse - so I’ll leave it here as a placeholder. See you at Mastodon!”
@@
Figure 5: Cross-platform engagement mapping of Twitter and Mastodon; counts of scholars classified by Twitter status (rows) and Mastodon usage (columns).
@@
We cross-mapped Twitter activity levels with Mastodon role categories, as shown in Figure 5. Our analysis reveals that 39.95% of the sampled scholars are relatively active on both Mastodon (long-term or persistent users) and Twitter (active users). Among those who are inactive on Twitter or have locked their accounts, 56.14% are long-term or persistent users of Mastodon. However, this group of dedicated Mastodon users represents only 7.44% of the total scholars included in the cross-platform analysis. This suggests that while there is a core group of loyal users who have fully migrated to Mastodon, they make up a relatively small portion of the broader group of early academic adopters.
@@
We also plotted the follower and following counts for the Mastodon and Twitter accounts of the same 3,131 scholars and the absolute differences in the metrics, as shown in Figure A4. The difference in connection scale is clear: while scholars who migrated to Mastodon made efforts to establish a new community, these networks remain far smaller than the connections many of them had built on Twitter over the years. Additionally, we plotted the Twitter account creation date against the total number of tweets, displayed in Figure A5. The color of the points represents the total number of likes received by each account. A great proportion of scholars created their Twitter accounts between 2008 and 2012, indicating that the deep, long-standing networks, high levels of engagement, and rich posting histories developed on Twitter are not easily replicable or discarded. This entrenched presence highlights the difficulty of transitioning to a new platform like Mastodon without sacrificing years of accumulated influence and community ties.
RQ2: Modeling Long-Term Retention
In previous sections, we have identified multiple factors that could potentially influence persistence on Mastodon. To understand the effect of these factors on persistence of Mastodon usage, we model them using a Cox proportional hazards model (Cox 1972). We selected this model to account for right censored data (users remaining active beyond the observation window) and to estimate the hazard ratios of covariates without assuming baseline hazard distribution.
Mastodon Survival Analysis
We used a Mastodon-centric model as our primary analysis and a cross-platform extension to understand the roles of Mastodon- and Twitter-derived features. The covariates used can be grouped as follows:
Mastodon activity metrics. These represent users’ platform activity level upon migration and social network metrics: the initial number of followers, following, and posts.
Federated interaction diversity. To capture how broadly a user engages across the Fediverse beyond their home server, we compute a network diversity metric from interactions. For each user, we extract the set of outgoing interactions and map each interaction to a target server. We then compute Shannon entropy (Lin 2002) over the distribution of target servers, yielding higher values when a user’s outgoing interactions are spread across many servers rather than concentrated within a single server. Our primary model uses diversity measured over the first 30 days of observed activity and we report sensitivity checks using alternative definitions (14 days, 60 days, and the first 50 outgoing interactions) in Figure A6 and A7.
Account and server-level features. We capture users’ profile settings, the network structure of their server, and the server type. Specifically, we consider whether the account is locked (binary), indicating if their posts are private and/or discoverable (binary), which indicates whether a user opted into being indexed to appear in search. At the server level, we identify the server type (binary) as either field-/topic-specific or general purpose by manually examining their public descriptions. A complete list of servers is provided in Appendix A4. At the server level, we define in-degree ratio as the proportion of inbound connections relative to total connections and out-degree ratio as the proportion of outbound connections relative to total connections. Additionally, we categorize users based on their academic discipline, creating a binary indicator for Multidisciplinary scholars in Table A1 to test if these communities show a distinct retention patterns.
Cross-platform metrics. These capture users’ premigration history on Twitter. We consider the followers and following counts on Twitter, the total number of Tweets since account creation, and the age of their Twitter account in days.
In both models, for each account, the number of days between our first data collection (11-23-2022) and the account’s last recorded posting activity within the study window was used as survival time. We treat a user having experienced the event of becoming inactive if their last posting occurs before the end of our data collection (10-27-2023). The counts covariates are log transformed and z-score normalized to handle the covariate distribution and comparable coefficients. We used a penalization ($L_{2}=0.1$) to prevent overfitting.
Mastodon-centric model
Our first model used the complete dataset of user accounts and features extracted exclusively from Mastodon data. We filtered out accounts on isolated servers and those with incomplete data, reducing our sample from 7,542 to a final account count of 7,357 users. The model achieved a strong predictive capability (concordance of 0.69) with several factors significantly predicting the risk of becoming inactive on Mastodon. The initial number of a user’s followers ($HR=0.72,p<0.0005$) and number of posts ($HR=0.74,p<0.001$) were strongest protective factors; one standard deviation increase in either was associated with around 30% lower likelihood of becoming inactive (refer to Appendix Figure A6). The initial number of accounts followed by a user did not have a significant effect ($HR=1.02,p=0.262$).
For the server type and profile settings, being on a topic-specific server was a strong protective factor ($HR=0.84,p<0.0005$), aligning with our observations from the information security server that tailored communities are associated with higher user retention. Notably, users identified as Multidisciplinary scholars exhibited higher retention rates with a hazard ratio of $0.59$ $(p<0.001)$. Furthermore, having a discoverable profile was also significantly linked to lower attrition ($HR=0.93,p=0.009$), suggesting that being able to be found and connected is linked to persistent usage. In contrast, having the profile locked was not found to be a statistically significant predictor. The out-degree ratio was a risk factor ($HR=1.80,p<0.001$) suggesting servers that are heavily outward looking are associated with lower user persistence. However, the proportion of inbound connections was not a significant predictor.
Federated interaction diversity is a robust protective predictor. In the 30-day diversity model, network diversity is associated with lower attrition ($HR=0.84,p<0.001$), indicating that a one standard deviation increase in early cross-server interaction diversity corresponds to an approximately 16% lower hazard of becoming inactive. Sensitivity checks using alternative windows (14 days, 60 days, and first 50 outgoing interactions) yield consistent results (Appendix Table X), with effect sizes ranging from $HR=0.89$ (14 days) to $HR=0.75$ (first 50 outgoing interactions).
Cross-platform model
Our second model was fitted using the subset of 3,131 users for whom we had complete cross-platform features from both Mastodon and Twitter. This allowed us to understand how a user’s Twitter activity related to their retention on Mastodon. The cross-platform model achieved concordance of 0.69 indicating that addition of Twitter features does not add much overall predictive power. The addition of Twitter covariates provides a granular view of retention correlates. Notably, the association with interaction network diversity remains robust ($HR=0.86,p<0.001$), indicating that cross-server engagement predicts retention independent of a user’s prior status on Twitter.
Among covariates used, the association of metrics extracted from Mastodon remained largely consistent with the first model (see Figure 6). Initial followers on the platform ($HR=0.72,p<0.0005$) and number of posts ($HR=0.77,p<0.001$) were still strong factors linked to retention. Similarly, server type ($HR=0.79,p<0.001$) and discoverability ($HR=0.86,p<0.001$) remained significant and positively associated with persistence. However, unlike the Mastodon-centric model, server structural metrics and Multidisciplinary scholar indicator were no longer significant predictors of attrition in the cross platform model.
Results of modeling covariates from Twitter revealed nuanced effects of some of the cross-platform user characteristics on persistence. A large number of Twitter following ($HR=1.13,p<0.0005$) and followers ($HR=1.11,p<0.0005$) were significant risk factors for becoming inactive on Mastodon. In contrast, a higher number of Twitter posts ($HR=0.92,p<0.001$) and older Twitter account age ($HR=0.88,p<0.0005$) were linked to higher retention.
The results from the cross-platform analysis shows a clear relationship with Twitter history. A large Twitter network is associated with higher attrition, while a long history of Twitter activity is linked to persistence showing divergent associations for these factors. Users with a well-established and high posting volume on Twitter were more committed to migration but the size of their network may have been too difficult to replicate on Mastodon pulling them back to Twitter. This suggests that a successful migration of a community from a well-established platform like Twitter requires not only individual commitment but also a critical mass of connections that can migrate to the new platform.
Figure 6: Plot showing hazard ratio and 95% CI for each covariate in the cross-platform model. Vertical dashed line at HR of 1.0 represents no effect.
Contextual Factors and Platform Environment
The model identifies measurable predictors of sustained use, yet migration and retention dynamics also unfold within a broader contextual environment that is not fully observable from activity traces alone. In this section, we examine three such contextual factors: the emergence of competing alternative platforms, public discourse surrounding the academic migration, and Mastodon’s platform affordances. These factors help contextualize the conditions under which early adopters navigated migration and sustained participation.
Alternatives to Mastodon
@@
In addition to Mastodon, other centralized alternatives to Twitter have emerged in recent years, and many academics explored these venues. Unlike Mastodon, these platforms provided familiar interfaces and ease of transitioning. Those present on these alternative platforms often added their account information in their Twitter account descriptions to make it easier for others to discover and connect with them on these new platforms. To understand the extent to which academics were exploring the alternatives,we analyzed their usernames and user descriptions for mentions of these platforms. Specifically, we focused on mentions of Bluesky [^9] https://bsky.app/ and Threads [^10] https://www.threads.net/ using keyword matching as well as user name pattern matching. For Mastodon and Bluesky, we also included scholars who use decentralized platform-style usernames as their Twitter usernames (e.g., <<username>>@<<servername>>).
@@
Figure 7: UpSet plot summarizing explicit cross-platform self-identification; intersection bars indicate counts of scholars mentioning Mastodon, Bluesky, and Threads (alone or jointly) in their profiles.
Public Discourse on Academic Migration
To examine public discourse surrounding early academic migration, we analyze public posts from Twitter and Mastodon that explicitly reference academic migration, collected using a predefined keyword-based retrieval strategy (see Section Data). We plot the daily frequency of these posts and engagement counts over time aggregated by week in Figure 8. We observe similar patterns on both platforms: discussions about the migration in the academic context experienced a noticeable spike during the first two months of the movement on both platforms. After the migration period, the conversations were less frequent on both platforms, they received more attention (favorites) on Twitter, indicating that the audience and connections of academics concerned about Mastodon have largely remained active on Twitter.
Qualitative inspection of highly engaged posts during peak periods reveals recurring themes related to uncertainty about platform stability and long-term viability. For example, Mastodon struggled to accommodate the abrupt influx of users. On the peak day of discussion, November 18, 2022:
“I have had a Mastodon presence for several months but haven’t used it. I still have a vain hope Twitter will sort itself out, but if not, I am <<scholar’s Mastodon account>>.” –Twitter Post
@@
Figure 8: Dynamic frequency of Twitter and Mastodon conversations related to academics migration and corresponding engagement metrics (public interaction and favorite counts), aggregated by week.
@@
“My humanities folks have largely migrated to mastodon, but my science and public health feeds are still here on Twitter. More’s the pity.” –Twitter Post
Despite these challenges, public efforts to encourage a shift toward Mastodon continued throughout 2023. For example, one scholar highlighted its advantages in a June 2023 post on Twitter:
“Reasons to join us in the fediverse. 500 character limit. free edit button. Lots more #openaccess resources for #teachers, researchers, science communicators, lifelong learners. #STEM #education #openscience #K12 #HigherEd” –Twitter Post
Momentum was also visible within Mastodon itself. A widely circulated post from October 2023, which received 613 reblogs, signaled a clear departure from Twitter:
Farewell, [bird emoji]! After 14 years on Twitter/X, we feel it’s time to part ways. Recent developments there do not align with our core values. We stand for trustworthy information provision and are committed to open science. Follow us here for open science-related updates and news on our services and support for the academic community.[heart emoji] #OpenScience #AcademicLibrary #GoodbyeTwitter #ResearchSupport #EducationSupport –Masotdon Post
Amongst the discussions on Twitter, we again observe concurrent mentions of other platforms such as Bluesky and Threads. These have diluted the migration as users explore various platforms for different purposes. For instance,
“Hey if anyone’s still here on science Twitter and wondering where a lot of people have gone, they’re trying out bluesky and liking it. This time it’s real (not like mastodon). #academic” –Twitter Post
Taken together, these posts illustrate a mix of optimism about Mastodon’s potential and recurring practical concerns users have faced during the migration.
Mastodon’s Technical Affordances
Beyond the individual- and community-level factors identified above, structural characteristics of Mastodon, such as its moderation policies and discoverability mechanisms, can be considered additional factors contributing to the limited success of migration. In line with the network structure, moderation policy is also decentralized, as described [^11] https://docs.joinmastodon.org/admin/moderation/:
Moderation in Mastodon is always applied locally, i.e. as seen from the particular server. An admin or moderator on one server cannot affect a user on another server, they can only affect the local copy on their own server.
This decentralized moderation stands in sharp contrast to Twitter’s centralized enforcement system. While local control empowers individual communities to define and enforce their own standards, it also creates inconsistencies across the network. For users accustomed to a uniform, centralized policy environment such as Twitter’s, these variations may generate uncertainty or limit trust in the platform.
Discoverability presents another technical challenge as shown in the models. Unlike Twitter, which supports platform-wide search and algorithmic amplification of content, Mastodon limits search to hashtags and profile information  [^12] https://docs.joinmastodon.org/user/network/#search. As a result, newcomers may struggle to discover communities, topics, or influential users, creating barriers to community formation and sustained engagement.
Taken together, these limiting affordances could help explain why Mastodon, despite its alignment with values of autonomy and community governance, has struggled to attract and retain large numbers of migrating users from Twitter.
Importantly, these technical affordances are not experienced uniformly across the Fediverse. Because each Mastodon server operates with its own administrative history, governance structure, and moderation capacity, users’ experiences may vary substantially depending on the server they join. Servers established prior to the migration period may benefit from more mature moderation practices or greater administrative experience, which could influence user satisfaction and retention.
Discussion and Conclusions
@@
Results from longitudinal analysis showed that the initial surge in sign-ups did not translate into sustained long-term user engagement. This pattern is consistent with prior studies on Twitter to Mastodon migration in general user groups showing early migration waves often do not translate to stable long term activity with many users maintaining cross platform presence and continue using Twitter as a primary platform for engagement rather than fully switching (Jeong et al. 2023; He et al. 2023) and with broader findings that the transition away from Twitter remains incomplete for general audiences (Peña-Fernández et al. 2025). Most recent user activity on Twitter revealed that many academics in our dataset returned to using Twitter. Further analysis suggested that the established history and strong communities established on Twitter, with some over a decade, proved too significant to overcome. While prior studies observed this general return t Twitter (Jeong et al. 2023; He et al. 2023), our cross-platform analysis with Twitter features, suggest that for established scholars, the difficulty of replicating existing networks is correlated with attrition. Without a complete migration of user history and established community, professionals and other specialized groups may naturally gravitate back to familiar platforms where their communities are already robust.
Likewise, several factors were linked to the failure of Mastodon to retain users. First, its decentralized structure creates a steeper learning curve compared to Twitter. Competition from more user-friendly platforms like Bluesky and Threads perhaps drew users away from Mastodon. Beyond direct Twitter substitutes, professional platforms such as LinkedIn continue to serve distinct functions centered around academic communication (Utz and Breuer 2019), which may further diffuse scholars’ attention across platforms rather than concentrating migration on a single alternative. Our analysis adds to the understanding of challenges in migration of professional communities by identifying interaction diversity as a robust predictor of retention. Additionally, those who continued to post consistently on Mastodon did not enjoy the same level of engagement as on Twitter. That being said, select groups of academics stayed committed to Mastodon, making this migration less of a straightforward failure and more of a complex transition. Prior work suggests that social influence and a shared sense of identity or belonging can shape collective migration behavior (Cava et al. 2023). In our data, higher persistence among scholars on field or topic specific servers is consistent with the possibility of community specific servers supporting sustained engagement. This commitment is likely reinforced by self selection. For example, the most motivated early adapters may have actively sought out community specific servers.